# CMM-Math: A Chinese Multimodal Math Dataset To Evaluate and Enhance the Mathematics Reasoning of Large Multimodal Models

### v2

@@ Conclusion
In this paper, we introduced the CMM-Math dataset, a comprehensive Chinese multimodal mathematical dataset designed to evaluate and enhance the performance of large multimodal models (LMMs) in mathematical reasoning. CMM-Math is distinguished by its scale, diversity, and complexity, comprising over 28,000 high-quality samples that span 12 grade levels and include various problem types. This dataset serves as both a benchmark and a training resource, addressing the gap in non-English, specifically Chinese, multimodal mathematical datasets. Our experiments reveal that existing state-of-the-art LMMs struggle with the challenges posed by the CMM-Math dataset, highlighting the need for further advancements in this field. We also propose a new math-specific LMM, Math-LMM, which is trained through a three-stage process: foundational pre-training, foundational fine-tuning, and mathematical fine-tuning. The results from our evaluations demonstrate that Math-LMM significantly improves performance in multimodal mathematical reasoning with a relatively compact model size of approximately 8 billion parameters.

### v3

@@ 6. Conclusion
In this paper, we introduce CMM-Math, a comprehensive Chinese multimodal mathematical dataset designed to evaluate and enhance the performance of LMMs in mathematical reasoning. CMM-Math is distinguished by its scale, diversity, and complexity, comprising over 28,000 high-quality samples that span 12 grade levels and include various problem types. This dataset serves as both a benchmark and a training resource, addressing the gap in non-English, specifically Chinese, multimodal mathematical datasets. Our experiments reveal that existing state-of-the-art LMMs struggle with the challenges posed by the CMM-Math dataset, highlighting the need for further advancements in this field. We also propose a new math-specific LMM, Math-LMM, which is trained through a three-stage process: foundational pre-training, foundational fine-tuning, and mathematical fine-tuning. The results from our evaluations demonstrate that Math-LMM improves performance in multimodal mathematical reasoning by comparing with open-source LMMs.